OpenStudy (anonymous):

Find all value(s) of x for which ........ Find all value(s) of x for which ........ @Mathematics

OpenStudy (anonymous):

\[2x^{2}-ix+1=x^{2}+2ix+3\]

OpenStudy (anonymous):

I got \[\frac{3i +-i \sqrt{17}}{2}\]

OpenStudy (anonymous):

im getting i and 2i as an answer. \[2x^2-ix+1=x^2+2ix+3\iff x^2-3ix-2=0\iff (x-i)(x-2i)=0\]

OpenStudy (anonymous):

ahh what did I do haha